# Universal properties of the many-body Lanczos algorithm at finite size

### v1

@@ Abstract
Introduction —
Understanding many-body quantum dynamics is an outstanding challenge: predicting whether and how a setup will eventually thermalize is a problem of exceptional theoretical interest with ample practical applications in the context of synthetic quantum matter. Several theoretical frameworks can describe the late-time behavior of a high-temperature quantum many-body system: from the eigenstate thermalization hypothesis [1, 2] to the conventional [3, 4] and generalized hydrodynamics [5, 6]; from numerical techniques based on entanglement [7] or dissipation [8] to the Lanczos algorithm (LA) [9, 10].
In the last years, the LA has experienced a revival of interest due to the conjecture that the Lanczos coefficients satisfy a universal behaviour associated to the chaotic or integrable behaviour of the model [11]. Among the other applications, the LA is a useful tool also for understanding the dynamics of the edge modes of different spin chains [12, 13, 14, 15, 16], to study Floquet systems [17, 18, 19, 20], to probe state and operator complexity in holography [21, 22, 23, 24, 25] and many-body quantum systems [26, 27, 28, 29, 30, 31, 32, 33]. The possibility of inferring emerging hydrodynamic effects [34, 35, 36, 10, 37, 38, 39, 40, 41, 42] or late-time quantum coherences from these coefficients has been debated; recent studies have begun to provide affirmative indications [43, 44]. Yet, these approaches suffer from the fact that the Lanczos coefficients that can be computed numerically are soon affected by finite-size effects, whereas the aforementioned studies focused entirely on infinite-system behaviors: this puts a strong constraint on the possibility of utilizing these conjectures in practical numerical simulations.
In this letter, we show that the Lanczos coefficients display universal properties even when studied at finite size. We focus on the asymptotic infinite-time value of the autocorrelation function of a local observable at finite size and discuss its relation with the LA. We conjecture that in non-integrable systems these coefficients are strongly constrained by the late-time hydrodynamic tail in a way that we make precise. If the autocorrelation function under study is related to a strong or approximate zero-mode, other behaviors are conjectured. We support our analysis with numerical simulations. This work opens the path to the use of the LA in finite systems to extract information on the universal late-time properties of quantum many-body systems.
Preliminaries —
@@
where the operator $\mathcal{O}(t)$ is time-evolved in the Heisenberg picture and obeys $\partial_{t}\mathcal{O}(t)=i[H,\mathcal{O}(t)]$. We explicitly state the dependence on the system size $L$ in Eq. (1) because it plays a crucial role in the forthcoming discussions.
@@
with $b_{n}>0$ known as Lanczos coefficients. Specifically, an orthonormal basis of operators (with respect to the product $(\mathcal{O},\mathcal{O}^{\prime}):=\langle\mathcal{O}^{\dagger}\mathcal{O}^{\prime}\rangle$) is constructed from the Gram–Schmidt orthogonalization algorithm applied to the nested commutators between $H$ and $\mathcal{O}$ (details are given in Ref. [11]). In the following, we will focus on observables with a vanishing expectation value ($\langle\mathcal{O}\rangle=0$), unit norm ($\langle\mathcal{O}^{\dagger}\mathcal{O}\rangle=1$), and, unless specified, on one-dimensional $d=1$ systems with short-range interactions.
Figure 1: Lanczos coefficients computed for model (3) for several sizes, from $L=9$ to $L=12$, with periodic boundary conditions. The observable is $\sigma^{z}_{1}$ and the parameters are $[J,h_{x},h_{z}]=[1,1,1.5]$. The plot shows clearly the presence of an infinite-size behaviour for $n<n^{*}$ followed by a fluctuating plateau region that is affected by finite-size effects.
@@
$H=\sum_{j}\left(J\sigma_{j}^{x}\sigma_{j+1}^{x}+h_{x}\sigma_{j}^{x}+h_{z}\sigma_{j}^{z}\right).$ (3)
In Fig. 1 we show the Lanczos coefficients for the operator $\sigma^{z}_{1}$ computed for finite chains up to $L=12$ with periodic boundary conditions. The plot displays the initial universal behavior independent of $L$, which persists in the thermodynamic limit, and the subsequent plateau, which is a finite-size effect. The crossover can be intuitively understood because the application of subsequent commutators of $H$ progressively increases the support of $\mathcal{O}$ and indeed takes place at $n^{*}$ of the order of $L$. Similar properties are also found for the other operators that we will consider, $\sigma^{y}_{1}$ and $\sigma^{z}_{1}\sigma^{z}_{2}$.
Whereas so far much of the studies have focused on the universal properties of the Lanczos coefficients for $n<n^{*}$, current numerical techniques can reliably compute only a few tens of them, so that large $n$ scalings are numerically difficult to address. The goal of this work is to identify the universal scalings that characterize the coefficients for $n>n^{*}$, which are easily numerically accessible (in this study, we can consider several hundreds of them).
The conjecture —
We focus on the late-time limit of $C_{L}(\infty)=\lim_{t\to\infty}C_{L}(t)$, and in particular on its behavior as a function of $L$. We begin by stating a formula, derived in the End Matter (EM), that relates exactly $C_{L}(\infty)$ to the Lanczos coefficients $b_{n}$:
@@
This formula is amenable to a simple interpretation: the Lanczos Hamiltonian (2) has a zero-energy eigenvector localised at the boundary of the fictitious Lanczos chain; the denominator of (4) is simply its squared norm. We can conclude that (i) $C_{L}(\infty)$ is different from zero when such mode exists and is normalizable, and that (ii) $C_{L}(\infty)=0$ when it does not exist, namely its norm blows to infinity.
The convergence of the infinite sum at the denominator of (4) depends eventually on the large-$n$ behavior of the (squared) ratios between odd and even Lanczos coefficients. We introduce the notation
@@
In generic closed many-body systems the energy is conserved and the associated hydrodynamic transport always takes place. As a consequence, the autocorrelators of a local operator $\mathcal{O}$ decay generically algebrically in time as $\sim t^{-\nu}$ reaching eventually a plateau of order $C_{L}(\infty)\sim L^{-md}$; the integer number $m$ is identified by the first power of the Hamiltonian $H$ that overlaps with $\mathcal{O}$, satisfying $\langle\mathcal{O}H^{m}\rangle_{c}\neq 0$, where $\langle\dots\rangle_{c}$ denotes the connected cumulant. This mechanism has been elucidated in Ref. [45], and a conjecture, named relaxation-overlap inequality, has been proposed: it predicts $\nu\leq dm/z$, with $z$ the dynamical critical exponent.
The behavior of the Lanczos coefficients for $n<n^{*}$ in models displaying a hydrodynamic behavior has been deeply investigated. Generically, these models display at leading order $b_{n}\sim n$ (for $d=1$ the scaling is $b_{n}\sim n/\ln{n}$ ) [11]); the subleading staggering proportional to $(-1)^{n}$ has been recently elucidated in Refs. [43, 44], and it is directly associated with the algebraic decay $\sim t^{-\nu}$.
@@
In particular, $\Gamma_{n}(L)$ should be independent of $n$ and positive, while consequently the ratio between odd and even Lanczos coefficients should be smaller than $1$. This prediction is expected to be true on average, meaning that fluctuations around this leading trend are possible and indeed found in the numerical data. Technical details relating our conjecture (6) to the predictions for $n<n^{*}$ are discussed in the EM.
@@
Even if transport is present, local operators do not show a hydrodynamic behavior when they do not overlap with conserved charges: this can happen, for example, in the presence of discrete symmetries like time-reversal symmetry or spatial inversion, where odd observables have vanishing overlaps with even conserved charges. In these situations, $C_{L}(\infty)$ vanishes or is exponentially small in system size, and we conjecture
@@
Specifically, it is possible either that $\Gamma_{n}$ is positive but converges to 00 sufficiently fast, or that it becomes negative at large $n$; in both cases, the denominator of (4) diverges and the plateau vanishes.
@@
Recently, situations where $C_{L}(\infty)$ is nonzero and independent of system size have attracted considerable attention. This is what happens in the presence of conserved local operators, known as strong zero modes [13, 16], where the system retains memory of the initial state even in the thermodynamic limit. A paradigmatic quantum system displaying a strong zero mode is the Su-Schrieffer-Heeger model (SSH) [46, 47] in its topological phase. For boundary fermionic operators, the LA produces a sequence of $b_{n}$ that oscillate perfectly between two values and terminates after $n\sim L$ [11].
Additionally, in interacting systems researchers have found approximate zero modes, characterized by a long-lived plateau of the autocorrelation function whose value does not depend on $L$ and that drops to zero after a given long time [13, 16]. There, for $n<n^{*}\sim L$, the Lanczos coefficients that have been numerically analyzed scale as $b_{n}\sim n+(-1)^{n}c$, with $c>0$, combining a linear growth in $n$ (as a consequence of non-integrability) to a superimposed staggering. For $n>n^{*}$ finite-size effects (described in the next section) kick in.
@@
asymptotically in $n$. This inequality guarantees mathematically the existence of a strong zero mode: when the zero modes are approximate, violations are expected for sufficiently large $n$.
Analytical arguments —
@@
in other words, we focus on those cases where either $\Gamma_{n}$ decays to zero algebraically in $n$ or converges to a (positive or negative) constant. By studying the behaviour of the cumulative products (details are reported in the EM), we find
@@
$C_{L}(\infty)\neq 0\;\Leftrightarrow\;\begin{cases}0\leq\beta<1\text{ and }\alpha>0,\\ \beta=1\text{ and }\alpha>2.\\ \end{cases}$ (11)
We now proceed with an estimate of the scaling with $L$ of $C_{L}(\infty)$. Assuming the leading trend $b_{n}\sim n$ in the region $n<n^{*}$ that was conjectured in Ref. [11] for generic systems, the first part of the series that appears at the denominator of Eq. (4) behaves as (see EM for details)
@@
In order to estimate the remaining part of the series, we first identify the scaling (proven in the EM)
@@
The Conjecture 2 concerns those cases where the denominator in Eq. (4) diverges. The formula in Eq. (7) is proposed to guarantee either a fast approach to 00 of the rate $\Gamma_{n}$ or a convergence to a negative value. By comparing it with the list of the converging cases given in Eq. (11), we see that it encompasses all non-converging situations.
We now consider Conjecture 3. As we have already discussed, in the presence of strong zero modes, the $b_{n}$ oscillate between two values and drop to zero after $n\sim L$. For small $n$, it is thus true that $\Gamma_{n}$ is constant. The infinite series at the denominator of Eq. (4) should become become a finite sum, whose convergence is automatically assured. This scenario is compatible with the conjectured inequality (8) because we are essentially saying that $\Gamma_{n}=+\infty$ for $n$ larger than the termination value.
@@
Numerics —
We now present some numerical results to support the three conjectures. The numerical study is performed with an exact-diagonalization algorithm using up to 600 GB of RAM; since the computation of the $b_{n}$ coefficients with the LA is numerically unstable, we thus resort to the full Gram-Schmidt orthogonalization scheme. The latter comes at a significant numerical cost: whereas with the LA we could compute several thousands of $b_{n}$ for setups up to $14$ sites, here we can only present data for $12$ or $13$ sites with several hundreds of coefficients. It is a remarkable fact discussed in the Supplementary Material (SM) that the coefficients computed with the standard LA algorithm describe the same autocorrelation function and satisfy the same universal scaling laws that we present here below.
We focus on the model in Eq. (3) and consider the three observables $\mathcal{O}=\sigma^{z}_{1}$, $\sigma^{z}_{1}\sigma^{z}_{2}$ and $\sigma^{y}_{1}$. According to the analysis presented in Ref. [45], $\sigma^{z}_{1}$ overlaps (in the infinite temperature state) with $H$ and $\sigma^{z}_{1}\sigma^{z}_{2}$ with $H^{2}$: as a consequence, the plateaus $C_{L}(\infty)$ scale respectively as $L^{-1}$ and $L^{-2}$. Hence, according to the Conjecture 1, $\Gamma_{n}(L)$ is expected to approach a positive constant $\sim L^{-(m+1)}$, with $m=1,2$ respectively.
Figure 2: The cumulative product $F(n)$ for the operators $\sigma^{z}_{1}$ (left panel) and $\sigma^{z}_{1}\sigma^{z}_{2}$ (right panel). The data refer to the Ising Hamiltonian (3) with parameters $[J,h_{x},h_{z}]=[1,1,1.5]$, and the dashed black lines show the fit with bi-exponential curves.
Figure 3: The cumulative products in Fig. 3 for the operator $\sigma^{z}_{1}$ ($m=1$) rescaled as $-\log F(n)\cdot L^{m+1}/n$. Assuming conjecture (6), the curves are expected to collapse for large $L$.
$F(n):=\prod^{n}_{n^{\prime}=0}e^{-\Gamma_{n^{*}+2n^{\prime}}(L)}.$ (15)
For both operators, the data are consistent with the presence of a decay rate that decreases with the system size. In order to check whether the scaling of the rate with $L$ is compatible with Eq. (6), in Fig. 3 we plot the quantity $-\log F(n)\cdot L^{m+1}/n$ as a function of $n$ for different sizes: these curves are expected to collapse for large $L$ toward a constant value. The results for $\sigma^{z}_{1}$ in Fig. 3 display a good agreement with the conjecture; the data for $\sigma^{z}_{1}\sigma^{z}_{2}$ are less convincing, and a better agreement would be found if larger values of $n$ were accessible: we comment the issue in the SM.
We now move to the study of the operator $\sigma^{y}_{1}$, which does not overlap with any power of $H$ (being odd under time reversal) [45]; the infinite-time plateau value of $C_{L}(\infty)$ is expected to vanish and thus it allows to test Conjecture 2 in Eq. (7). Our numerics finds that $\Gamma_{n}(L)$ fluctuates around a negative value and that $F(n)$ grows as a function of $n$, as shown in Fig. 4 (left panel).
Finally, we consider the Hamiltonian
with open boundary conditions and parameters $\mu=0.2$, $U=0.5$; the model is known to host a boundary zero mode at zero temperature that becomes approximate when the temperature is increased [16]. We probe the operator $\mathcal{O}=\sigma^{x}_{1}$ at the left boundary of the chain in the infinite temperature state. We show the results in Fig. 4 (right panel): the data for $F(n)$ are compatible with a saturation at large $n$, consistent with $\Gamma_{n}\simeq 0$, implying $C_{L}(\infty)=0$ according to Eq. (11). We remark that, in contrast, if the zero mode was exact, we would have observed an exponential decay of $F(n)$ with a size-independent decay rate $\Gamma$. By comparing the two panels of Fig. 4, we observe a marked qualitative difference between the two cases, which share the common property $C_{L}(\infty)=0$, but are associated to two qualitatively distinct physical situations (absence of hydrodynamic projection versus approximate zero mode).
Conclusions and outlook —
We have considered the LA for quantum many-body systems, and we have focused on the properties of the Lanczos coefficients $b_{n}$, computing several hundreds of them for finite-size systems. By focusing on the universal properties of autocorrelation functions in generic models, we have presented a threefold conjecture on the scaling of the ratios $(b_{n}/b_{n+1})^{2}$, with $n$ odd, in the large $n$ limit. The main goal of our study is to show that the finite-size behavior of the Lanczos coefficients contains universal information that can be studied with state-of-the-art numerics.
This work opens intriguing research directions. A natural question is to understand the reason behind the bi-exponential behavior found in Fig. 3. Moreover, after having studied the late-time plateau of $C_{L}(t)$, it is natural to focus on the algebraic decay $t^{-\nu}$ (partially addressed in Ref. [44]), and to investigate how the finite-size Lanczos coefficients give rise to such a behavior. Also, the precise mechanism by which the finite-size crossover of $C_{L}(t)$, from decay to plateau, occurs at the level of the LA is currently unclear. Finally, our study of the Lanczos coefficients has highlighted an erratic behavior in the region $n>n^{*}$; whether the latter has some universal properties, and it can be studied as a stochastic process, is a question that deserves to be addressed.
Acknowledgements.
Acknowledgments
End Matter
Derivation of Eq. (4) for $C_{L}(\infty)$
First, we briefly summarize the relation between the spectrum of $\mathcal{L}$ and the autocorrelation function $C_{L}(t)$: this is standard material, and we refer the reader to Ref. [11] for details. Using the Lanczos basis, the first basis element |$\ket{1}$ corresponds to the seed operator $\mathcal{O}$ and the autocorrelation function can be written as
$C_{L}(t)=\bra{1}e^{i\mathcal{L}t}\ket{1}.$ (17)
In particular, its Fourier transform coincides with the projection of the resolvent operator of $\mathcal{L}$ onto |$\ket{1}$:
[…]
with
$\Phi(\omega):=\frac{1}{\pi}\text{Im}\bra{1}\frac{1}{\omega-i0^{+}-\mathcal{L}}\ket{1}$ (19)
[…]
Thus, we try to solve |$\mathcal{L}\ket{\phi}=0$, with |$\ket{\phi}$ a column vector with components $\{\phi_{n}\}_{n>0}$ and $\mathcal{L}$ the matrix in Eq. (2). This is performed iteratively, since the matrix $\mathcal{L}$ is tridiagonal, and the condition is equivalent to
As a result, $\phi_{n}$ vanishes for even $n$, while it can be written efficiently in terms of $\phi_{n-2}$ for $n$ odd as $\phi_{n}=-(b_{n-2}/b_{n-1})\phi_{n-2}$. Putting everything together, and requiring that the state |$\ket{\phi}$ is properly normalized, that is $\sum_{n}|\phi_{n}|^{2}=1$, one finds
Putting everything together, and studying the small $\omega$ behavior of $\Phi(\omega)$, we find
$\Phi(\omega)=\frac{|\innerproduct{1}{\phi}|^{2}}{\pi}\text{Im}\frac{1}{\omega-i0^{+}}+\dots=|\phi_{1}|^{2}\delta(\omega)+\dots,$ (23)
Lastly, we note a technical but important point. So far, we have implicitly treated $\mathcal{L}$ as if it was a semi-infinite matrix, which potentially can have both discrete and real spectrum. On the other hand, whenever $L$ is finite, that is the relevant case of our discussion, the Hilbert space of operators is finite, although exponentially large in the system size, and the spectrum is always discrete. As a consequence, the Lanczos matrix $\mathcal{L}$ is always finite and the sequence of ${b_{n}}$ always ends at $n\sim e^{O(L)}$. Nonetheless, the information coming from the last Lanczos coefficients only affects the dynamics of the model at time scales that are exponentially large in the system sizes: these times, where revivals occur due to Poincaré recurrences, are physically irrelevant. In contrast, finite-size effects at order $t\sim L^{z}$, associated with transport, with dynamical exponent $z$, are much more accessible; therefore, in practice, we can regard $\mathcal{L}$ as if it were a semi-infinite matrix as long as these time-scales are concerned.
Analytical arguments: technical details —

### v2

@@ Abstract
I Introduction
Understanding many-body quantum dynamics is an outstanding challenge: predicting whether and how a setup will eventually thermalize is a problem of exceptional theoretical interest with ample practical applications in the context of synthetic quantum matter. So far, several theoretical frameworks can be used to address the late-time behavior of a high-temperature quantum many-body system: from the eigenstate thermalization hypothesis [1, 2] to the conventional [3, 4] and generalized hydrodynamics [5, 6]; from numerical techniques based on entanglement [7] or dissipation [8] to the Lanczos algorithm (LA) [9, 10].
In the last years, the LA has experienced a revival of interest due to the conjecture that the Lanczos coefficients satisfy a universal behaviour associated to the chaotic or integrable behaviour of the model [11]. Among the other applications, the LA is a useful tool also for understanding the dynamics of the edge modes of different spin chains [12, 13, 14, 15, 16], to study Floquet systems [17, 18, 19, 20, 21], to probe state and operator complexity in holography [22, 23, 24, 25, 26] and many-body quantum systems [27, 28, 29, 30, 31, 32, 33, 34]. The possibility of inferring emerging hydrodynamic effects [35, 36, 37, 10, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47] has been also discussed. Yet, these approaches suffer from the fact that the Lanczos coefficients that can be computed numerically are soon affected by finite-size effects, whereas the aforementioned studies focused entirely on infinite-system behaviors: this puts a strong constraint on the possibility of utilizing these conjectures in practical numerical simulations.
In this work, we show that the Lanczos coefficients display universal properties even when studied at finite size. We extend the Universal Operator Growth Hypothesis [11], expected to be valid for generic many-body systems in infinite lattices; we find that the Lanczos coefficients follow general finite-size patterns that are largely independent from the microscopic details of the interactions. We obtain this result by focusing on the asymptotic infinite-time value of the autocorrelation function of a local observable at finite size and by discussing its relation with the LA. We present a threefold conjecture that covers several situations. We support our analysis and our claim of universality with numerical simulations on several lattice models with qualitatively different properties. This work opens the path to the use of the LA in finite systems to extract information on the universal late-time properties of quantum many-body systems.
We structure the manuscript as follows: in Sec. II, we introduce the theoretical framework and summarize the Lanczos algorithm, focusing on its application to autocorrelation functions and the infinite-time behavior of local observables. In Sec. III, we present our central conjecture, which connects the late-time plateau of autocorrelation functions to the scaling of Lanczos coefficients, and discuss its implications for hydrodynamic behavior, vanishing plateaus, and strong zero modes. In Sec. IV, we provide analytical arguments to support our conjecture, while Sec. V is dedicated to numerical benchmarks across various models, including long-range and higher-dimensional systems. We discuss conclusions and outlook in Sec. VI and address technical details of the numerics in Appendix A.
II Preliminaries
@@
where the operator $\mathcal{O}(t)$ is time-evolved in the Heisenberg picture and obeys $\partial_{t}\mathcal{O}(t)=i[H,\mathcal{O}(t)]$. In Eq. (1) we explicitly state the dependence on $L$ because it plays a key role in the forthcoming discussions.
@@
with $b_{n}>0$ known as Lanczos coefficients. Specifically, an orthonormal basis of operators (with respect to the product $(\mathcal{O},\mathcal{O}^{\prime}):=\langle\mathcal{O}^{\dagger}\mathcal{O}^{\prime}\rangle$) is constructed from the Gram–Schmidt orthogonalization algorithm applied to the nested commutators between $H$ and $\mathcal{O}$ (see Ref. [11]). We will focus on observables with a vanishing expectation value ($\langle\mathcal{O}\rangle=0$) and unit norm ($\langle\mathcal{O}^{\dagger}\mathcal{O}\rangle=1$).
Figure 1: Left: Lanczos coefficients computed for model (3) for several sizes, from $L=10$ to $L=13$, with periodic boundary conditions. The observable is $\sigma^{z}_{1}$ and the parameters are $J,h_{x},h_{z}]=[1,1,1.5]$. The plot clearly shows the presence of an infinite-size behaviour for $n<n^{*}$ followed by a fluctuating plateau region affected by finite-size effects. Right: Corresponding correlation functions $C_{L}(t)$ as a function of time $t$; a size-dependent plateau is observed at late times.
@@
$H=\sum_{j}\left(J\sigma_{j}^{x}\sigma_{j+1}^{x}+h_{x}\sigma_{j}^{x}+h_{z}\sigma_{j}^{z}\right),$ (3)
where $\sigma^{\alpha}_{j}$, $\alpha=x,y,z$, are the Pauli operators at the $j$-th site. In the left panel of Fig. 1 we show the Lanczos coefficients for the operator $\sigma^{z}_{1}$ computed for finite chains up to $L=13$ with periodic boundary conditions. The plot displays the initial universal behavior independent of $L$, which persists in the thermodynamic limit, and the subsequent plateau, which is a finite-size effect. The crossover can be intuitively understood because the application of subsequent commutators of $H$ progressively increases the support of $\mathcal{O}$ and indeed takes place at $n^{*}$ of the order of $L$. Similar properties are also found for the other operators that we will consider, $\sigma^{y}_{1}$ and $\sigma^{z}_{1}\sigma^{z}_{2}$. As a comparison, we plot the corresponding correlation functions $C_{L}(t)$ in the right panel of Fig. 1: these (approximately) coincide for $t$ sufficiently small, and then show a plateau depending on the system size $L$.
Whereas so far many of the studies have focused on the universal properties of the Lanczos coefficients for $n<n^{*}$, current numerical techniques can reliably compute only a few tens of them, so that large $n$ scalings are numerically difficult to address. The goal of this work is to identify the universal scalings that characterize the coefficients for $n>n^{*}$. Specifically, we want to understand the relation between the plateau of the autocorrelation functions and the corresponding Lanczos coefficients.
III The conjecture
We focus on the late-time limit of $C_{L}(\infty)=\lim_{t\to\infty}C_{L}(t)$, and in particular on its behavior as a function of $L$. We begin by stating a formula, derived in Sec. III.1 (see also Ref. [44]), that relates exactly $C_{L}(\infty)$ to the Lanczos coefficients $b_{n}$:
@@ III The conjecture
This formula is amenable to a simple interpretation: the Lanczos Hamiltonian (2) has a zero-energy eigenvector localised at the boundary of the fictitious Lanczos chain; the denominator of (4) is simply its squared norm. We can conclude that (i) $C_{L}(\infty)$ is different from zero when such a mode exists and is normalizable, and that (ii) $C_{L}(\infty)=0$ when it does not exist, namely its norm blows to infinity.
The convergence of the infinite sum at the denominator of (4) ultimately depends on the large-$n$ behavior of the (squared) ratios between odd and even Lanczos coefficients. We introduce the notation
@@
In generic closed many-body systems the energy is conserved and the associated hydrodynamic transport always takes place. As a consequence, the autocorrelators of a local operator $\mathcal{O}$ decay generically algebrically in time as $\sim t^{-\nu}$ reaching eventually a plateau of order $C_{L}(\infty)\sim L^{-md}$; the integer number $m$ is identified by the first power of the Hamiltonian $H$ that overlaps with $\mathcal{O}$, satisfying $\langle\mathcal{O}H^{m}\rangle_{c}\neq 0$, where $\langle\dots\rangle_{c}$ denotes the connected cumulant. This mechanism was elucidated in Ref. [48], where a conjecture, named relaxation-overlap inequality, has been proposed: it predicts $\nu\leq dm/z$, with $z$ the dynamical critical exponent.
[…]
@@
In particular, $\Gamma_{n}(L)$ should be independent of $n$ and positive, while consequently the ratio between odd and even Lanczos coefficients should be smaller than $1$. This prediction is expected to be true on average, meaning that fluctuations around this leading trend are possible and indeed found in the numerical data. Technical details relating our conjecture (6) to the predictions for $n<n^{*}$ are discussed in Sec. IV.1.
@@
Even if transport is present, local operators do not show a hydrodynamic behavior when they do not overlap with conserved charges. In these situations, $C_{L}(\infty)$ vanishes or is exponentially small in system size, and we conjecture
@@
Specifically, it is possible either that $\Gamma_{n}$ is positive but converges to 0 sufficiently fast, or that $\Gamma_{n}$ becomes negative at large $n$; in both cases, the denominator of (4) diverges and the plateau vanishes.
@@
Recently, situations where $C_{L}(\infty)$ is nonzero and independent of system size have attracted considerable attention. This is what happens in the presence of conserved local operators, known as strong zero modes [13, 16], where the system retains memory of the initial state even in the thermodynamic limit (see e.g. [49, 50]). For boundary fermionic operators, the LA produces a sequence of $b_{n}$ that oscillate perfectly between two values and terminates after $n\sim L$ [11].
Additionally, in interacting systems researchers have found approximate zero modes, characterized by a long-lived plateau of the autocorrelation function whose value does not depend on $L$ and that drops to zero after a given long time [13, 16]. There, for $n<n^{*}\sim L$, the Lanczos coefficients that have been numerically analyzed scale as $b_{n}\sim n+(-1)^{n}c$, with $c>0$. For $n>n^{*}$ finite-size effects, described in the next section, kick in.
@@
asymptotically in $n$. This inequality guarantees mathematically the existence of a strong zero mode: when the zero modes are approximate, violations are expected.
III.1 Derivation of Eq. (4) for $C_{L}(\infty)$
[…]
$C_{L}(t)=\int^{\infty}_{-\infty}d\omega\ \Phi(\omega)e^{i\omega t},$ (10)
with
[…]
$\begin{split}C_{L}(\infty):=\lim_{T\rightarrow\infty}\frac{1}{T}\int^{T/2}_{-T/2}dt\ C_{L}(t),\\ \Phi(\omega)=C_{L}(\infty)\delta(\omega)+\dots,\end{split}$ (12)
[…]
By studying the small $\omega$ behavior of $\Phi(\omega)$, we find
$\Phi(\omega)=\frac{|\langle 1\mid\phi\rangle|^{2}}{\pi}\text{Im}\frac{1}{\omega-i0^{+}}+\dots=|\phi_{1}|^{2}\delta(\omega)+\dots,$ (15)
Lastly, we note a technical but important point. So far, we have implicitly treated $\mathcal{L}$ as if it were a semi-infinite matrix, which potentially can have both discrete and real spectrum. On the other hand, whenever $L$ is finite, that is the relevant case of our discussion, the Hilbert space of operators is finite, although exponentially large in the system size, and the spectrum is always discrete. As a consequence, the Lanczos matrix $\mathcal{L}$ is always finite and the sequence of $b_{n}$} always ends at $n\sim e^{O(L)}$. Nonetheless, the information coming from the last Lanczos coefficients only affects the dynamics of the model at time scales that are exponentially large in the system sizes: these times, where revivals occur due to Poincaré recurrences, are physically irrelevant. In contrast, finite-size effects at order $t\sim L^{z}$, associated with transport, with dynamical exponent $z$, are much more accessible; therefore, in practice, we can regard $\mathcal{L}$ as if it were a semi-infinite matrix as long as these time-scales are concerned.
IV Analytical arguments supporting the conjecture
@@
in other words, we focus on those cases where either $\Gamma_{n}$ decays to zero algebraically in $n$ or converges to a (positive or negative) constant. By studying the behaviour of the cumulative products (details are reported in Sec. IV.1), we find
@@
[…]
We now proceed with an estimate of the scaling of $C_{L}(\infty)$ with $L$. Assuming the leading trend $b_{n}\sim n$ in the region $n<n^{*}$ that was conjectured in Ref. [11] for generic systems, the first part of the series that appears at the denominator of Eq. (4) behaves as (see Sec. IV.1 for details)
@@
In order to estimate the remaining part of the series, we first identify the scaling (proven in the IV.1)
@@
The Conjecture 2 concerns those cases where the denominator in Eq. (4) diverges. The formula in Eq. (7) is proposed to guarantee either a fast approach to 0 of the rate $\Gamma_{n}$ or a convergence to a negative value. By comparing it with the list of the converging cases given in Eq. (18), we see that it encompasses all non-converging situations.
We now consider Conjecture 3. As we have already discussed, in the presence of strong zero modes, the $b_{n}$ oscillate between two values and drop to zero after $n\sim L$. For small $n$, it is thus true that $\Gamma_{n}$ is constant. The infinite series at the denominator of Eq. (4) should become a finite sum, whose convergence is automatically assured. This scenario is compatible with the conjectured inequality (8) because we are essentially saying that $\Gamma_{n}=+\infty$ for $n$ larger than the termination value.
@@
IV.1 Technical details